Estimate the Joule-Thomson coefficient of steam at (i) 3 MPa and 300°°C and (ii) 6 MPa and 500°C.


1
Expert's answer
2021-04-21T07:05:53-0400

"T_{3.5MPa} = \\300+ \dfrac{ 2994.3-2978.4}{h(350°C) -h(300°C)}(50)"


"300 +\dfrac{2994.3 -2978.4}{3104.9-2978.4}(50) = 306.3°C"


"T_{2.5MPa} = \\300 - \dfrac{h(300°C)- 2994.3}{h(300°C) -h(250°C)}(50)"


"300- \dfrac{3009.6-2994.3}{3009.6-2880.9}(50) = 294.1°C"